# SySLLM: Generating Synthesized Policy Summaries for Reinforcement Learning Agents Using Large Language Models

### v1

@@ Abstract
Policies generated by Reinforcement Learning (RL) algorithms can be difficult to describe to users, as they result from the interplay between complex reward structures and neural network-based representations. This combination often leads to unpredictable behaviors, making policies challenging to analyze and posing significant obstacles to fostering human trust in real-world applications. Global policy summarization methods aim to describe agent behavior through a demonstration of actions in a subset of world-states. However, users can only watch a limited number of demonstrations, restricting their understanding of policies. Moreover, those methods overly rely on user interpretation, as they do not synthesize observations into coherent patterns. In this work, we present SySLLM (Synthesized Summary using LLMs), a novel method that employs synthesis summarization, utilizing large language models’ (LLMs) extensive world knowledge and ability to capture patterns, to generate textual summaries of policies. Specifically, an expert evaluation demonstrates that the proposed approach generates summaries that capture the main insights generated by experts while not resulting in significant hallucinations. Additionally, a user study shows that SySLLM summaries are preferred over demonstration-based policy summaries and match or surpass their performance in objective agent identification tasks.  [^1] All relevant resources for this study, such as SySLLM summaries, HIGHLIGHTS summaries, expert summaries, experience videos, the source code, and other materials are available at https://osf.io/c78jb/?view_only=76fbf59131d4409bac222e5929a654f2.
@@ 1 Introduction
Explainable Reinforcement Learning (XRL) methods aim to enable human users to understand agents’ policies and decision-making processes. Such understanding is crucial for fostering trust and ensuring the real-world deployment of RL agents. Several approaches to XRL have been developed in recent years [25], including saliency maps that visualize the agent’s attention, reward bars that break down expected utility into components, and causal explanations that clarify why actions were taken. Other methods extract logical rules to describe when actions are taken, but these approaches rely on predefined logical representations and often focus on individual actions rather than broader behavior patterns. Despite these advancements, experiments reveal that human users still struggle to form accurate mental models of agent behavior [6, 19, 23].
Recent advancements in large, pre-trained language models (LLMs) present a novel opportunity to enhance explanation methods for reinforcement learning (RL) policies. These models exhibit exceptional in-context learning abilities and possess extensive commonsense knowledge about human behavior, which makes them promising tools for generating meaningful, language-based explanations of RL policies. LLMs have already been employed in constructing world models and planning with domain-independent planners, showcasing their potential for reasoning and providing explanation [14].
@@
In this paper, we aim to leverage LLMs to generate comprehensive global explanations of RL agent policies. Specifically, our goal is to use LLMs to produce natural language summaries that encapsulate key patterns of behavior, capabilities, and limitations of RL policies. This approach seeks to build upon previous work that utilized demonstration-based policy summaries [4], where agent policies are described through selected trajectories in video. However, these methods are inherently constrained by the representativeness of the selected scenarios and the limited number of demonstrations that can feasibly be observed. They do not synthesize observations into coherent summaries, leaving users to rely heavily on their own interpretations to discern patterns. In contrast, we hypothesize that LLM-generated summaries can effectively distill these patterns into concise, coherent narratives, providing users with a clearer and more accessible understanding of agent policies.
Leveraging LLMs to generate accurate and comprehensible policy summaries is a non-trivial task. While LLMs exhibit remarkable reasoning abilities, they are prone to hallucinations—producing outputs that seem plausible but deviate from the agent’s true behavior [16]. This raises questions about whether such descriptions can reliably reflect the nuances of RL agent policies. Moreover, their utility heavily depends on end users’ ability to interpret them. Laypeople, who often lack familiarity with RL concepts, may struggle to form accurate mental models of agent behavior. Natural language explanations, while accessible, risk oversimplifying or failing to convey subtle patterns [6].
A further challenge lies in capturing the spatial and temporal relationships inherent in RL environments. These relationships evolve continuously over many steps and episodes: spatial dependencies include changing object positions, while temporal dependencies arise from sequential actions. Since LLMs are trained primarily on static text, they may struggle to model these dynamics effectively [24]. Together, these factors complicate the task of distilling and communicating agent behaviors in a way that remains faithful to the policy while being accessible to users.
Our method, Synthesized Summary using LLMs (SySLLM), leverages LLMs to generate comprehensive summaries of RL agent behavior. The proposed pipeline first collects the agent’s interactions within the environment (Figure 1) and converts these into natural language descriptions of observations and actions. These descriptions are then used to prompt an LLM, which synthesizes a natural language summary that captures key patterns in the agent’s behavior (Figure 2). We implement this methodology across five distinct environments, generating language-based summaries for agents with diverse behavioral patterns.
To evaluate SySLLM, we employ two complementary methods. First, we compare its summaries to those created by six RL experts, finding substantial agreement between the two. Second, we conduct a user study where participants interact with two types of policy summaries: visual summaries generated by the HIGHLIGHTS-DIV algorithm [3] and language-based summaries generated by SySLLM. Participants rated SySLLM higher in subjective preferences and performed equally well or better at identifying agents based on these summaries. These results highlight the effectiveness of SySLLM in both comprehensibility and practical utility.
The main contributions of our work are:
  • We introduce Synthesized Summary using LLMs (SySLLM), a novel and versatile method that generates holistic, synthesized, language-based summaries of RL agent policies using LLMs.
  • We validate the correctness of SySLLM by comparing its summaries to those of RL experts, showing substantial alignment and demonstrating that SySLLM accurately captures key policy details.
@@
  • We conduct a large-scale user study, demonstrating that participants prefer SySLLM over a state-of-the-art demonstration-based policy summarization method in a summary preference task and perform equally well or better in an agent identification task.
@@ 2 Background and Related Work
Explainable RL (XRL)
Our work extends the body of research concerned with explaining RL policies and particularly the line of work on global policy explanations (see Amitai and Amir [5] for a recent survey). Policy summarization approaches present users with demonstrations of the agent acting in the environment to convey its behavior. Several criteria for extracting such summaries have been proposed in the literature, including state importance, diversity, visitation frequency, and policy reconstruction [3, 17, 23, 18, 34]. While these approaches were shown to improve people’s ability to assess agent capabilities and predict agent behavior, users can only grasp a limited amount of demonstrations, and experiments show that their interpretation of policies highly depends on the subset of examples shown to them [23].
Other works propose learning a simplified model of the policy (e.g., a decision tree or a set of logical rules) that captures the essence of the agent’s decision-making [38, 8, 7]. Yet, these approaches are also limited due to the tradeoff between the accuracy of the simplified model and its interpretability to users.
We hypothesize that LLMs may offer a complementary approach to global explanations by capturing meaningful patterns of agent behavior and presenting them in understandable language to users.
Pre-trained Language Models in RL
Large Language Models (LLMs) are probabilistic text models trained on substantial text datasets. Models like GPT [30], which are autoregressive, excel in text generation by sequentially predicting future words based on the preceding text. Such models can be useful in the context of RL as they encode common-sense knowledge about the world. Conversely, models such as BERT [12], which are based on the masked framework, focus on understanding deep sentence semantics by training to recognize parts of text that are intentionally obscured. Encoder-decoder models merge these traits to adeptly handle tasks requiring both the interpretation and creation of text. These models have been extensively pre-trained on large-scale text corpora, equipping them with the ability to tackle complex language processing tasks with minimal specific instruction, often achieving impressive results [9]. This extensive training allows them to efficiently process and summarize complex narratives, and to identify relevant patterns within large datasets.
Recent work has developed methods that use LLMs in the training of RL policies. For example, LLMs are being used to guide intrinsic exploration by suggesting plausibly useful goals and providing intrinsic rewards [13, 27]. Other works use LLMs to generate useful state abstractions [29, 28]. Additionally, LLMs are employed to provide natural language instructions and dynamic feedback, enhancing agent learning in complex, open-ended environments [39]. Differently from these approaches, in this work, we aim to use LLMs to explain RL policies.
3 Problem Statement
We define a partially observed Markov Decision Process (POMDP) as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,\gamma\rangle$, where observations $o\in\mathcal{O}$ are derived from states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$ via $O(o|s,a)$. The dynamics are governed by the transition function $T(s^{\prime}|s,a)$, with $R$ as the reward function and $\gamma$ as the discount factor. POMDPs generalize Markov Decision Processes (MDPs), where observations fully reveal the state.
The agent operates under a stochastic policy $\pi:\mathcal{O}\rightarrow\Delta(\mathcal{A})$, mapping each observation $o\in\mathcal{O}$ to a probability distribution over actions $\mathcal{A}$. The objective of the policy is to maximize the expected cumulative discounted reward.
Given a POMDP environment and the agent’s policy, our goal is to generate a natural language summary that encapsulates the agent’s strategy and decision-making process. This summary explains the policy’s operational dynamics, effectiveness, and behavior across scenarios. Formally, a summarization is a mapping from $(\mathcal{M},\Pi)$ to $\mathcal{T}$, where $(\mathcal{M},\Pi)$ represents environment-policy pairs, and $\mathcal{T}$ is the set of natural language summaries.
4 Methodology
@@
Our method consists of several key components, which we formalize in Algorithm 1. In the first phase (illustrated in Figure 1), we run the agent in its environment and perform a captioning procedure to convert experiences into natural language. These captioned experiences are logged into an experience dataset (Sections 4.1 and 4.2). In the second phase, which we illustrate in Figure 2, we take the experience dataset as input and output a natural language summary that aims to capture patterns in the agent’s policy (Section 4.3). To do so, we compose a prompt with the formatted experience dataset that is then sent to an LLM to generate the summary.
@@ 4.2 Collecting the Agent’s Experience Dataset
We describe the first phase of Algorithm 1, detailed in Lines 114.
We assume that the environment ($ENV$) supports two functions. The first function is RESET, which resets the environment to its initial state and returns the initial observation, and STEP, which takes an action as input and returns the next observation, the reward for that action, a boolean indicating whether the episode has ended, and additional information.
First, we initialize the experience dataset $ED$ (Line 1). The for loop (Line 2) iterates over multiple episodes, initializing the step counter $t=1$ at the beginning of each episode (Line 3) and obtaining the initial observation $o_{1}$ (Line 4). The variables $epReward$ and $done$ are also initialized to track the cumulative reward and episode completion status, respectively (Lines 56). The while loop (Line 7) continues until the episode concludes.
During each iteration of the while loop, an action is sampled from the policy $\pi$ (Line 8), and both the observation and action are converted into natural language representations by the captioners. These representations are added to the experience dataset along with the cumulative reward (Line 9). Subsequently, the environment is updated using the selected action, and the next observation is retrieved (Line 10).
Finally, the cumulative reward and step index are updated accordingly (Lines 1112). The information stored in the experience dataset is detailed in Section B of the Appendix.

### v2

# From Actions to Words: Towards Abstractive-Textual Policy Summarization in RL

@@ Abstract
Policies generated by Reinforcement Learning (RL) algorithms are difficult to explain to users, as they emerge from the interaction of complex reward structures and neural network representations. Consequently, analyzing and predicting agent behavior can be challenging, undermining user trust in real-world applications. To facilitate user understanding, current methods for global policy summarization typically rely on videos that demonstrate agent behavior in a subset of world states. However, users can only watch a limited number of demonstrations, constraining their understanding. Moreover, these methods place the burden of interpretation on users by presenting raw behaviors rather than synthesizing them into coherent patterns. To resolve these issues, we introduce SySLLM (Synthesized Summary using Large Language Models), advocating for a new paradigm of abstractive-textual policy explanations. By leveraging Large Language Models (LLMs)—which possess extensive world knowledge and pattern synthesis capabilities—SySLLM generates textual summaries that provide structured and comprehensible explanations of agent policies. SySLLM demonstrates that LLMs can interpret spatio-temporally structured descriptions of state-action trajectories from an RL agent and generate valuable policy insights in a zero-shot setting, without any prior knowledge or fine-tuning. Our evaluation shows that SySLLM captures key insights, such as goal preferences and exploration strategies, that were also identified by human experts. Furthermore, in a large-scale user study (with 200 participants), SySLLM summaries were preferred over demonstration-based summaries (HIGHLIGHTS) by a clear majority (75.5%) of participants.
journal: Artificial Intelligence
[…]
@@ 1 Introduction
Reinforcement learning (RL) agents are increasingly deployed in high-stakes domains, yet their opaque decision-making processes remain a barrier to trust and adoption. Explainable RL (XRL) methods aim to bridge this gap by surfacing interpretable accounts of agent behavior. Despite this promise, current XRL methods face significant limitations. Saliency maps [1, 2, 3, 4] highlight critical states but offload pattern discovery to users. Trajectory demonstrations [5, 6, 7, 8] constrain the understanding to cherry-picked scenarios. Rule extraction methods [9, 10, 11, 12], while providing structured representations, often struggle to capture the full complexity of learned policies. Many of these approaches share a common challenge: they present transformed observations rather than synthesized insights, requiring users to manually extrapolate agent intent, adaptability, and failure modes from fragmented evidence [13]. Recent advances in large language models (LLMs) offer a transformative alternative. With their capacity for abstraction, commonsense reasoning, and natural language generation [14], LLMs hold unique potential to synthesize agent trajectories into human-interpretable narratives. However, this promise comes with a challenge, LLMs’ static text-based training contrasts with the dynamic, spatiotemporal nature of reinforcement learning environments [15]. Early applications of LLMs in RL, such as world modeling [16], hint at their promise, but generating faithful global policy summaries—cohesive accounts that generalize beyond isolated states—remains unexplored.
@@
In this work, we advocate for an abstractive-textual approach for policy summarization. We introduce SySLLM, a framework that leverages LLMs to generate holistic summaries of RL policies by systematically distilling agent trajectories into structured narratives. SySLLM addresses the limitations of previous methods through a two-stage process: (1) converting raw agent-environment interactions into temporally grounded language descriptions (Fig. 1), and (2) prompting LLMs to synthesize these into coherent summaries that capture behavioral patterns, strategic tendencies, and environment-specific adaptations (Fig. 2). SySLLM performs impressively well—it captures agent preferences (as induced by different reward functions), identifies sub-optimal behavior in under-trained agents, and distinguishes between diverse exploration strategies. Beyond summarizing the agent’s behavior, SySLLM offers insights into the effectiveness of the agent’s strategy.
SySLLM demonstrates that LLMs can interpret spatiotemporally structured descriptions of state-action trajectories from an RL agent. While these descriptions are written in natural language, they differ significantly from the kinds of text LLMs are typically trained on, such as stories or dialogue. Instead, they represent a distinct modality: sequential, decision-driven behavior grounded in the dynamics of an environment. Despite this, SySLLM shows that LLMs can process such unfamiliar input and, in a zero-shot setting—without any prior knowledge or fine-tuning—generate valuable policy insights.
Our framework is validated through multi-faceted evaluation. First, comparisons with RL experts reveal that SySLLM summaries align closely with expert annotations, capturing nuanced behaviors that demonstration-based methods do not explicitly highlight. Second, a large-scale user study shows that participants prefer SySLLM over state-of-the-art visual summaries such as HIGHLIGHTS [17], while also achieving equal or better accuracy in agent identification tasks.
In summary, our main contributions are the following:
  • We introduce SySLLM, a framework that takes a step toward generating abstractive-textual explanations of RL policies using LLMs. SySLLM synthesizes structured narratives that reveal agent intent, adaptability, and limitations by systematically analyzing agent trajectories, enabling users to grasp these behaviors without manual pattern matching. SySLLM eases the burden of interpreting agent behavior by digesting complex dynamics into coherent summaries. To the best of our knowledge, this is the first work to produce global, abstractive explanations of an RL agent’s policy in natural language.
  • Generated summaries align with RL experts. We evaluate the effectiveness of SySLLM by comparing its summaries with those written by RL experts, demonstrating substantial alignment and accurate coverage of key policy details. To evaluate faithfulness—an area often prone to issues in automatic summarization [18, 19]—we additionally compute precision scores. Our results show that SySLLM maintains high precision overall, indicating that its summaries rarely introduce unsupported claims.
@@
  • Users prefer our textual summaries. We conduct a large-scale user study with 200 participants showing that participants clearly prefer SySLLM over a leading demonstration-based policy summarization method on a summary preference task (75.5%), while performing equally well or better in an agent identification task.
@@ 2 Background and Related Work
2.1 Explainable Reinforcement Learning
Explainable reinforcement learning (XRL) aims to make the behavior of RL agents more understandable to human users. Traditional XRL techniques fall into several major categories: saliency-based methods, demonstration-based summaries, surrogate models, and causal explanations [20].
Saliency and Visualization Techniques
Early XRL work often adopted visualization tools such as saliency maps [1, 2, 3, 4] to indicate which parts of the input space influenced the agent’s decisions. While these tools provide some insight, they require users to manually interpret the spatial and temporal relevance of highlighted regions, leading to potential misinterpretation. Studies show these methods are local and fragile [21], providing little insight into the long-term policy structure.
Demonstration-Based Summaries
Another common approach involves selecting representative episodes or critical states that demonstrate agent behavior [5, 6, 7, 8]. These methods allow users to observe behavior directly, but rely heavily on cherry-picked scenarios, with high cognitive load placed on the user to infer strategic intent from limited data. Work by [22] shows that users often struggle to generalize from such demonstrations, particularly when the agent behavior is stochastic.
Surrogate Models
An alternative strategy uses model distillation [23] to learn interpretable approximations of agent behavior, such as decision trees [24, 25]. These methods strike a balance between interpretability and fidelity: overly simple models fail to capture nuances of the agent’s strategy, while complex surrogates become as opaque as the original policy. Recent works explore bounding MDPs [26] or abstract policy graphs [27] to improve this trade-off, but scalability remains an issue.
Causal Explanations
A complementary direction models the agent’s reasoning through structural causal models or influence graphs [28, 29]. These allow users to query “why” an agent took a specific action and offer contrastive explanations that align with how humans seek understanding. However, such models often require significant prior knowledge of the domain or agent internals, limiting their applicability.
Reward Decomposition
Other methods explain agent behavior by breaking down the overall reward into meaningful components, such as safety, efficiency, or task-specific objectives. This allows explanations that clarify how each reward type influences the agent’s decisions and highlights trade-offs between competing objectives [30, 31, 32]. These approaches help users understand agent priorities and decision rationale by making the contribution of each reward component explicit.
2.2 Large Language Models in Reinforcement Learning
Pre-trained Large Language Models (LLMs) have revolutionized NLP through their ability to distill patterns from vast datasets into coherent text. Autoregressive models like GPT [33] generate human-like narratives by predicting token sequences, while encoder-decoder architectures excel at summarizing complex inputs [34]. Those recent advances have introduced new possibilities for interacting with and explaining RL agents. While much prior work integrates LLMs into the training process e.g., guiding exploration [35] or building world models [16], less attention has been paid to their potential as explanation generators. [36] explore using LLMs to generate local post-hoc explanations of individual decisions based on distilled behavior representations, but such approaches do not aim to provide global summaries of agent behavior.
Trajectory Annotation and Feedback
Some works use LLMs to narrate agent behavior in real-time or provide textual feedback. [37, 38] demonstrate that LLMs can be prompted to interpret actions or suggest improvements mid-episode. These studies show promise for LLM-based commentary, though they often lack grounding in actual agent dynamics, leading to inconsistencies.
Model-Based Reasoning and Simulation
In world modeling applications, LLMs have been used to generate or refine symbolic models of environment dynamics [39]. Here, the LLM acts as a high-level planner or simulator, encoding causal knowledge that guides behavior. This further validates the hypothesis that LLMs can represent abstract behavioral structures, which is essential for high-level summarization.
These applications largely treat LLMs as agents or environment models, not as explainers. When LLMs are used for explanation, they must be carefully grounded in observed experience to avoid plausible-sounding but inaccurate generalizations [40].
2.3 Natural Language Explanations for RL Agents
The use of natural language as a modality for RL explanations offers distinct advantages over visual or rule-based approaches. Language allows for abstraction, temporal coherence, and commonsense reasoning—all of which are challenging to express in saliency maps or decision trees.
Template-Based Explanations
Early works, such as [41], demonstrated that combining learned MDP structures with templated natural language generation could enable contrastive explanations (e.g., “Why didn’t you do X?”) through predicate-based state abstraction, allowing generalization across policy architectures. Later, simpler approaches reemerged, such as [10], which employed rigid, static templates to directly translate policies into interpretable statements (e.g., “If near wall, turn left”), prioritizing accessibility but inheriting brittleness and limited expressivity. More recent systems like [42] also rely on predefined action-condition mappings, facing similar limitations.
Rationalization and Free-Form Generation
Subsequent work adopted neural methods to generate textual rationales. [43] introduced AI rationalization training sequence-to-sequence models to translate trajectories into natural language. Later, [44] introduced graph-based policy summaries that enhance user understanding by clustering states and abstracting the state space.
While those natural language explanations methods offer interpretability, they are inherently limited by their reliance on handcrafted rules and rigid structures, making them brittle and unable to express nuanced or unexpected behaviors. Rationalization methods, which translate individual trajectories into free-form text, tend to focus on local justifications rather than providing a holistic view of agent behavior. These approaches often lack scalability across diverse scenarios and struggle to generalize patterns observed over many episodes. Moreover, they typically do not leverage broader world knowledge or incorporate high-level reasoning, which can limit the clarity, depth, and utility of the resulting explanations for end users.
3 Motivation for Abstractive-Textual Explanations
Despite recent advances in explainable reinforcement learning (XRL), existing methods face a persistent trade-off between expressiveness, scalability, and accessibility. Extractive approaches—such as saliency maps or curated demonstrations—highlight fragments of agent behavior but often fail to integrate them into coherent, high-level narratives. Symbolic models, including logical rules and decision trees, offer structural clarity but require handcrafted features and often struggle to scale in complex or high-dimensional settings.
These limitations point to a need for explanation methods that can synthesize global behavioral insights—capturing patterns, adaptation, and strategy over time—without relying on structured input representations or human-crafted abstractions.
We advocate for a textual-abstractive paradigm in XRL, where natural language is used to summarize agent behavior across episodes. Text offers a flexible medium for conveying temporal dependencies, recurring behaviors, and decision-making logic in a format accessible to non-experts. Unlike extractive or symbolic methods, textual explanations can operate over raw behavioral data without predefined rules or model internals, making them broadly applicable across domains.
Recent advances in large language models (LLMs) open new opportunities for explainability in reinforcement learning. With SySLLM, we introduce a textual-abstractive paradigm that uses natural language to summarize agent behavior over time—capturing patterns, strategic intent, and adaptation. In this work, we show that LLMs are particularly well-suited for this approach, when guided by an appropriate framework, they can recognize sequential structure and express it in fluent, context-aware language. We view this paradigm as a holistic and scalable complement to existing XRL methods, and as part of a broader shift toward language-based interfaces for interpreting learned behavior.
4 SySLLM Method
@@
Our SySLLM (Synthesized Summary using LLMs) method consists of several key components, which we formalize in Algorithm 1. In the first phase (illustrated in Figure 1), we run the agent in its environment and perform a captioning procedure to convert experiences into natural language. These captioned experiences are logged into an experience dataset (Sections 4.2 and 4.3). In the second phase, which we illustrate in Figure 2, we take the experience dataset as input and output a natural language summary that aims to capture patterns in the agent’s policy (Section 4.4). To do so, we compose a prompt with the formatted experience dataset that is then sent to an LLM to generate the summary.
4.1 Problem Statement
We define a partially observed Markov Decision Process (POMDP) as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,\gamma\rangle$, where observations $o\in\mathcal{O}$ are derived from states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$ via $O(o|s,a)$. The dynamics are governed by the transition function $T(s^{\prime}|s,a)$, with $R$ as the reward function and $\gamma$ as the discount factor. POMDPs generalize Markov Decision Processes (MDPs), where observations fully reveal the state. The agent operates under a stochastic policy $\pi:\mathcal{O}\rightarrow\Delta(\mathcal{A})$, mapping each observation $o\in\mathcal{O}$ to a probability distribution over actions $\mathcal{A}$. The objective of the policy is to maximize the expected cumulative discounted reward. Given a POMDP environment and the agent’s policy, our goal is to generate a natural language summary that encapsulates the agent’s strategy and decision-making process. This summary explains the policy’s operational dynamics, effectiveness, and behavior across scenarios. Formally, a summarization is a mapping from $(\mathcal{M},\Pi)$ to $\mathcal{T}$, where $(\mathcal{M},\Pi)$ represents environment-policy pairs, and $\mathcal{T}$ is the set of natural language summaries.
@@ 4.3 Collecting the Agent’s Experience Dataset
We describe the first phase of Algorithm 1, detailed in Lines 114. We assume that the environment ($ENV$) supports two functions. The first function is RESET, which resets the environment to its initial state and returns the initial observation, and STEP, which takes an action as input and returns the next observation, the reward for that action, a boolean indicating whether the episode has ended, and additional information. First, we initialize the experience dataset $ED$ (Line 1). The for loop (Line 2) iterates over multiple episodes, initializing the step counter $t=1$ at the beginning of each episode (Line 3) and obtaining the initial observation $o_{1}$ (Line 4). The variables $epReward$ and donedone are also initialized to track the cumulative reward and episode completion status, respectively (Lines 56). The while loop (Line 7) continues until the episode concludes. During each iteration of the while loop, an action is sampled from the policy $\pi$ (Line 8), and both the observation and action are converted into natural language representations by the captioners. These representations are added to the experience dataset along with the cumulative reward (Line 9). Subsequently, the environment is updated using the selected action, and the next observation is retrieved (Line 10). Finally, the cumulative reward and step index are updated accordingly (Lines 1112). The information stored in the experience dataset is detailed in B.